Convert (3–√,−1) into polar coordinates (r,θ) so that r≥0 and 0≤θ<2π.


1
Expert's answer
2022-06-03T02:53:34-0400
"r^2=(-\sqrt{3})^2+(-1)^2=4"





"r\ge0=>r=\sqrt{4}=2"

Quadrant III



"\tan \theta=\dfrac{-1}{-\sqrt{3}}=\dfrac{1}{\sqrt{3}}""\theta=\pi+\tan^{-1}(\dfrac{1}{\sqrt{3}})=\dfrac{7\pi}{6}"

"(2, \dfrac{7\pi}{6})"